# Directed Hypercube Routing, a Generalized Lehman-Ron Theorem, and Monotonicity Testing

### v1

@@ 4.1 From flows to directed isoperimetry
We start with a notion of the “directed volume” of a set.
@@
This brings us to a sort of “intellectual starting point” for this paper. The theorems above clearly show how directed isoperimetry and flows are intimately connected. Moreover, statements like Theorem 4.3 suggest relations between flows with edge capacities, and flow with vertex capacities. We were motivated to see if the KMS theorem (Theorem 4.4) could be proven from a flow perspective.
@@ 4.2 Connections between conjectures
The main upshot of this lemma is that one can “break up” the $(S,\overline{S})$ routing problem into a collection of $(S_{i},T_{i})$ routing problems, where the $S_{i},T_{i}$ are level subsets. Moreover, the interaction between the various cover graphs is limited, because of the last bullet point. Given that the separation distance of $S$ is $r$, we believe that for a constant fraction (by total size) of the matched pairs $(S_{i},T_{i};\phi_{i})$, the distance of these pairs is $\Omega(r)$. If Conjecture 1.3 is true, we can route $\Omega(r|S_{i}|)$ units of edge disjoint flow with vertex congestion $r$. Any vertex participates is at most $2r$ such flow. We had hoped to overlay such flows and get an overall vertex congestion of $O(r^{2})$. Unfortunately, a direct overlay of flows leads to an edge congestion of $O(r)$, which is not useful for Conjecture 4.10. Nonetheless, it felt that a proof of Conjecture 1.3 with the proof techniques of Theorem 4.8 might yield insight into Conjecture 4.10.

### v2

@@ 4.1 From flows to directed isoperimetry
We start with a notion of the ”directed volume” of a set.
@@
This brings us to a sort of ”intellectual starting point” for this paper. The theorems above clearly show how directed isoperimetry and flows are intimately connected. Moreover, statements like Theorem 4.3 suggest relations between flows with edge capacities, and flow with vertex capacities. We were motivated to see if the KMS theorem (Theorem 4.4) could be proven from a flow perspective.
@@
We believe that Conjecture 4.10 is stronger than Theorem 4.4, because it explicitly involves the separation distance of $S$.
As an aside, the connection between flows and directed isoperimetry resolves an open question in [FHH+14] (Pg 19, ”Routing on the hypercube”). It is actually a direct consequence of Theorem 4.2.
Theorem 4.12.
Let $(S,T,\phi)$ be a matched pair where $S$ and $T$ are disjoint. There exist $|S|$ monotone edge disjoint paths from $S$ to $T$.
Proof.
Consider the directed hypercube and take a mincut separating $S$ from $T$. Construct a Boolean function that assigns $1$ to the ”$S$-side”, and 00 to the ”$T$-side”. All remaining vertices can be assigned values such that they do not participate in any monotonicity violation. Note that these vertices cannot be on a directed path from $S$ to $T$. (Process vertices according to the partial order. For $x$, set $f(x)$ to be $\max_{y\prec x:f(y)\ \textrm{assigned}}f(y)$. If no $f(y)$ is assigned, set $f(x)=0$. Observe that if $f(x)$ is assigned value $1$, then $x$ must be greater than some point in $S$. This means that $x$ cannot be less than any point in $T$, and hence does not create monotonicity violations.)
This function has distance to monotonicity at least $|S|/2^{n}$. So by Theorem 4.2, there are at least $|S|$ edges which have value $(1,0)$. These are precisely cut edges, from the $S$-side to the $T$-side. Hence, the cut value is at least $|S|$. Set up a flow problem on the directed hypercube where every edge has unit capacity, vertices in $S$ are sources, and vertices in $T$ are sinks. By the maxflow-mincut theorem, there is a flow of value at least $|S|$. This flow gives edge-disjoint paths from S to T. ∎
@@ 4.2 Connections between conjectures
The main upshot of this lemma is that one can ”break up” the $(S,\overline{S})$ routing problem into a collection of $(S_{i},T_{i})$ routing problems, where the $S_{i},T_{i}$ are level subsets. Moreover, the interaction between the various cover graphs is limited, because of the last bullet point. Given that the separation distance of $S$ is $r$, we believe that for a constant fraction (by total size) of the matched pairs $(S_{i},T_{i};\phi_{i})$, the distance of these pairs is $\Omega(r)$. If Conjecture 1.3 is true, we can route $\Omega(r|S_{i}|)$ units of edge disjoint flow with vertex congestion $r$. Any vertex participates is at most $2r$ such flow. We had hoped to overlay such flows and get an overall vertex congestion of $O(r^{2})$. Unfortunately, a direct overlay of flows leads to an edge congestion of $O(r)$, which is not useful for Conjecture 4.10. Nonetheless, it felt that a proof of Conjecture 1.3 with the proof techniques of Theorem 4.8 might yield insight into Conjecture 4.10.